OpenStudy (anonymous):

Can someone teach me how to isolate P1 from L=10Log(Pi/P0) if 10 is the base of the logarithm. I know the answer but I want is the procedure.

OpenStudy (anonymous):

205 fans... seriously???? O.o

OpenStudy (amistre64):

What does P1 represent? I see a P0..but am unsure what you need help with

OpenStudy (amistre64):

......i guess so lol

OpenStudy (amistre64):

must be an accounting error ;)

OpenStudy (amistre64):

...not only am I a sandwich...but a SUPER sandwich....oh the humanity!!

OpenStudy (anonymous):

just isolate P1 from L=10log(P1/Po). I need the procedure

OpenStudy (anonymous):

you just gotta use log rules, a base to the log of it is the base itself: L=10log(blah) so 10^L=10^10log(blah) so 10^L=blah

OpenStudy (anonymous):

amistre, congratulations on making a superhero not a joke, shows that you have helped alot of people

OpenStudy (anonymous):

or the quotient rule... log(x/y)=log(x)-log(y)

OpenStudy (amistre64):

ohh...so instead if "Pi" its P1? just trying to get a sense of the problem but log rules say that division is subtraction log(a/b) = log(a) - log(b)

OpenStudy (amistre64):

thankx oktal :) but I am still leary of how much "help" I have been ;)

OpenStudy (anonymous):

you have been a GREAT help to me, and I am sure to MANY others

OpenStudy (anonymous):

thank you guys!

OpenStudy (anonymous):

did we answer the question?

OpenStudy (anonymous):

click the fan button :DD

OpenStudy (amistre64):

L = 10log(P1/Po) L = log(P1^10/Po^10) L = 10(log(P1) - log(P0)) (L/10) + log(P0) = log(P1) 10^(L/10 + log(Po)) = P1 10^(L/10) 10^log(Po) = P1 this is what I get to so far :)

OpenStudy (anonymous):

i win! lolololol

OpenStudy (amistre64):

:) arman won yay!!

OpenStudy (anonymous):

Yeah you perfectly answered my question guys. I was messing with something that now makes me laugh. Thanks again

OpenStudy (anonymous):

fan buttonnn <3

OpenStudy (anonymous):

wait... whad i win?

OpenStudy (amistre64):

gotta watch out for those "things that make you laugh"....theyll turn on ya :)

OpenStudy (amistre64):

arman won a new car!!!! and ....and..... a christmas tree airfreshner to hang from the mirror and obscure the fuzzy dice :)

OpenStudy (anonymous):

lol yo i gotta go to chinese class now, wo shi zhong guo xue sheng...

OpenStudy (amistre64):

Ni hen hao...